# Dynamics of states of infinite quantum systems as a cornerstone of the second law of thermodynamics

### v1

@@ Abstract
In his basic work [Di30], Dirac emphasized that the basic building blocks of quantum mechanics are states and observables, states being vectors in a (finite-dimensional) Hilbert space, and observables, self-adjoint operators acting on these states. In the present paper we shall see that it is imperative, in order to arrive at a precise statement of the second law of thermodynamics in the sense of Claudius, to generalize this notion of state to infinite dimensions, or, more precisely, to infinite number of degrees of freedom (ndof). In order to sharpen our intuition, it appears to be useful to undertake a preliminary investigation of the situation in connection with classical dynamical systems, which leads us to the probabilistic properties of (deterministic) dynamical systems [LM85]. Since generic dynamical systems are chaotic, we choose to look at one of the simplest chaotic systems, the so-called dyadic map [LM85].
@@
We now come to the subject of states and observables for the dyadic map and chaotic states. For an alternative treatment of states of classical systems, see Ruelle’s article [Ru67]. We begin with a definition:
@@ 2.4 Dynamical properties of states of infinite systems related to the modified second law
Our choice of the initial state $\omega_{0}$ as an equilibrium state may also be expected to be expressible in terms of some properties of the dynamics, viz., of the function $t\to\omega_{t}$. For $T>0$ this is the well-known KMS boundary condition, see [Hu72]. We shall restrict ourselves in section 3 to the case $T=0$, and the analogous ground state condition is also well-known ([Hu72], Theorem 3.2): there is a function $F(z)$, analytic for $Imz<0$, uniformly bounded for $Imz\leq 0$, continuous on the real axis, such that, for $z=t$ (real)
@@
$(\omega_{0}\circ\tau^{1})\circ\tau\neq\omega_{0}\circ\tau^{1}\mbox{ for }t>0$ (51)
@@
time-arrow theorem The preparation part of a system undergoing an adiabatic transformation as defined in  2.4 necessarily breaks time-reversal invariance, that is, there exists some scalar function of the states and observables of the system which is not ivariant under time-reversal.
@@
2.4.3 Quantum chaotic states. The Krylov mechanism of approach to equilibrium: the role of dynamics and interactions
@@ Acknowledgements
I should like to thank Elliott Lieb and Heide Narnhofer, who have strongly influenced the present paper, through various scientific exchanges, very heartily. I am also indebted to the late Derek Robinson, who insisted on the relevance of the property of affinity, which plays an important role in this paper. Finally, I should like to thank Michael Kiessling for his kind permission to include Fig.2 of his paper with Albert, which turned out, unexpectedly, to describe the chaotic quantum dynamics of the whole class of Dyson models.

### v2

@@ Abstract
In his basic work [Di30], Dirac emphasized that the basic building blocks of quantum mechanics are states and observables, states being vectors in a (finite-dimensional) Hilbert space, and observables, self-adjoint operators acting on these states. In the present paper we shall see that it is imperative, in order to arrive at a precise statement of the second law of thermodynamics in the sense of Clausius, to generalize this notion of state to infinite dimensions, or, more precisely, to infinite number of degrees of freedom (ndof). In order to sharpen our intuition, it appears to be useful to undertake a preliminary investigation of the situation in connection with classical dynamical systems, which leads us to the probabilistic properties of (deterministic) dynamical systems [LM85]. Since generic dynamical systems are chaotic, we choose to look at one of the simplest chaotic systems, the so-called dyadic map [LM85].
@@
We now come to the subject of states and observables for the dyadic map and chaotic states. For an alternative treatment of states of classical systems, see Ruelle’s article [Ru67]. See also the paper by Illner and Neunzehrt [IN87], in which the change between the two topologies as below is held responsible for the fact that in the limit of a process with time-reversal invariance in the Boltzmann equation a “time-arrow” is generated. In order to characterize these two topologies, we begin with a definition.
@@ 2.4 Dynamical properties of states of infinite systems related to the modified second law
Our choice of the initial state $\omega_{0}$ as an equilibrium state may also be expected to be expressible in terms of some properties of the dynamics, viz., of the function $t\to\omega_{t}$. For temperature $T>0$ this is the well-known KMS boundary condition, see [Hu72]. We shall restrict ourselves in section 3 to the case $T=0$, and the analogous ground state condition is also well-known ([Hu72], Theorem 3.2): there is a function $F(z)$, analytic for $Imz<0$, uniformly bounded for $Imz\leq 0$, continuous on the real axis, such that, for $z=t$ (real)
@@
$(\omega_{0}\circ\tau^{1})\circ\tau_{t}\neq\omega_{0}\circ\tau^{1}\mbox{ for }t>0$ (51)
@@
time-arrow theorem The preparation part of a system undergoing an adiabatic transformation as defined in  2.4 necessarily breaks time-reversal invariance, that is, there exists some scalar function of the states and observables of the system which is not invariant under time-reversal.
@@
2.4.3 Quantum chaotic states: the role of dynamics and the classical character
@@
The chaotic nature of the dynamics does vary strongly with certain parameters, particularly with the time. The reader should compare Fig.2 of the paper by Albert and Kiessling, reproduced in the present paper, with their Fig.1, for small times, which displays significantly less chaos. We may understand this phenomenon intuitively if we recall that chaotic dynamics is a generic aspect of classical mechanics (classical dynamical systems), and we expect that the classical limit is defined by a condition of the form
$\frac{\hbar}{ET}\ll 1$ (64)
where $E$ denotes an energy value in the typical range allowed by the system, and $T$ a corresponding typical time value: the quantity $\frac{\hbar}{ET}$ is of course dimensionless. In our case, we may take in the Dyson model $E$ as the maximum value in  (60) and $T$ as the time (see  (59)). In this sense, comparison between the two figures in [AK17] also provides strong graphical evidence of the dictum: the limit of large times is classical!
@@ Acknowledgements
I should like to thank Elliott Lieb and Heide Narnhofer, who have strongly influenced the present paper, through various scientific exchanges, very heartily. Heide also helped me to improve the final version with several important remarks. I am also indebted to the late Derek Robinson, who insisted on the relevance of the property of affinity, which plays an important role in this paper. Finally, I should like to thank Michael Kiessling for his kind permission to include Fig.2 of his paper with Albert, which turned out, unexpectedly, to describe the chaotic quantum dynamics of the whole class of Dyson models.